Find the slope of the line, if it is defined.
Through (7, -4) and (3, 8)

Answer by nerdybill(7384)

Find the slope of the line, if it is defined.
Through (7, -4) and (3, 8)
m (slope) = (y2-y1)/(x2-x1)
m (slope) = (8-(-4))/(3-7)
m (slope) = (8+4)/(3-7)
m (slope) = (12)/(-3)
m (slope) = -4